OpenStudy (anonymous):

The following set of scores was recorded for the results of a statistics exam: 95, 98, 68, 77, 92, 74, 67, 86, 72, 79, 58, 95, 64, 81, 47, 85, 92, 84, 85 Which of the histograms below best represents the data? select one

OpenStudy (gorv):

where is histogram ??

OpenStudy (anonymous):

hang on

OpenStudy (dan815):

http://prntscr.com/50w9pd

OpenStudy (dan815):

here i plotted it for u

OpenStudy (perl):

dan, what program did you use to make that histogram

OpenStudy (anonymous):

@kohai yo can you help me?

OpenStudy (anonymous):

@PaulaLovesSchool13 @iGreen Thank you!!

OpenStudy (anonymous):

Wait, I didn't help?

OpenStudy (anonymous):

Oh well I meant thank you for your futuristic help haha

OpenStudy (anonymous):

XD

OpenStudy (igreen):

Okay, look the options separate the values into groups of 6: 41 - 50 51 - 60 61 - 70 71 - 80 81 - 90 91 - 100 Values: 47,58,64,67,68,72,74,77,79,81,84,85,85,86,92,92,95,95,98 Based on this you can see: 41 - 50 = 1 value 51 - 60 = 1 value 61 - 70 = 3 values 71 - 80 = 4 values 81 - 90 = 5 values 91 - 100 = 5 values

OpenStudy (igreen):

Which graph fits that? @MarthaHearst

OpenStudy (igreen):

Therefore, the answer is B.. @MarthaHearst

OpenStudy (anonymous):

I get it thanks so much @iGreen

OpenStudy (igreen):

No problem!